A movie theatre charges $7.50 for adults and $4.50 for children. The receipts for one showing of a movie were $540. If 80 tickets were sold, find how many tickets of each type were sold.
Answer by mananth(16946)

A movie theatre charges $7.50 for adults and $4.50 for children. The receipts for one showing of a movie were $540. If 80 tickets were sold, find how many tickets of each type were sold.
$7.50 for adults Let number of adults = x
$4.50 for children. number of children =y
Receipts = $540
tickets sold = 80
..
1...x+y=80
2...7.50x+4.50y=540
Multiply eq 1 by 7.5 and subtract eq2
7.5x+7.5y-7.5x-4.5y=60
3y=60
y=20 the number of children
number of adults = 60